# FlashMD: long-stride, universal prediction of molecular dynamics

### v1

@@ Abstract
Molecular dynamics (MD) provides insights into atomic-scale processes by integrating over time the equations that describe the motion of atoms under the action of interatomic forces. Machine learning models have substantially accelerated MD by providing inexpensive predictions of the forces, but they remain constrained to minuscule time integration steps, which are required by the fast time scale of atomic motion. In this work, we propose FlashMD, a method to predict the evolution of positions and momenta over strides that are between one and two orders of magnitude longer than typical MD time steps. We incorporate considerations on the mathematical and physical properties of Hamiltonian dynamics in the architecture, generalize the approach to allow the simulation of any thermodynamic ensemble, and carefully assess the possible failure modes of such a long-stride MD approach. We validate FlashMD’s accuracy in reproducing equilibrium and time-dependent properties, using both system-specific and general-purpose models, extending the ability of MD simulation to reach the long time scales needed to model microscopic processes of high scientific and technological relevance.
@@ 2.3 ML modeling of molecular dynamics trajectories
Generative ML techniques have been applied to directly sample the thermodynamically accessible system configurations [48, 49, 50, 51], which is especially useful for biomolecular systems with slow conformational transitions. Such models allow cheap prediction of the thermodynamic properties that conventionally require long MD simulations, but they disregard the time-dependent behavior of the systems and hence are unsuitable for investigating the physicochemical phenomena that can only be explained via the system’s dynamics. More recently, the implicit transfer operator has been proposed to extend Boltzmann generators to also recover, with stochastic trajectory, a measure of long-time dynamical processes [52].
@@
This is the class of methods that most closely resembles our approach. The ML models take the positions and momenta of atoms of the system at time step as and deterministically predict their changes at , hence “directly propagating” the dynamics. In MDnet by Zheng et al. [61], the chemical system is described as a graph, with the edge features incorporating both the relative positions and momenta between the atoms. The model then predicts the changes in the positions and momenta for a fixed large time step . Very recently, Thiemann et al. have demonstrated TrajCast [62], an autoregressive equivariant network for direct MD prediction. Their framework has been shown to achieve good accuracy in reproducing trajectories for individual molecular or bulk systems at a specific thermodynamic state point. It is also worth mentioning the Equivariant Graph Neural Operator [63] (EGNO) approach, which predicts the evolution of the system at multiple times using equivariant temporal convolution in Fourier space, and the Graph Network-based Simulators [64, 65] (GNS), which have been developed for arbitrary particle-based systems without the chemical context.
@@ 3 The FlashMD framework
In this section, we explain the design choices made for “FlashMD”, our proposed framework for the direct learning and prediction of MD trajectories. Given that MD shares with interatomic potentials the properties of symmetry and the use of atomic geometries as inputs, we propose that FlashMD should be built on top of similar GNN architectures. In this work, we choose the Point-Edge Transformer [18] (PET), although any GNN architecture could be used. We make two physically motivated modifications to adapt it to the prediction of trajectories: (i) Each node state is also initialized using the particle momentum , encoded via a multi-layer perceptron, in addition to the chemical species of the atom under consideration. The initialization of the edge states remains unchanged, and it includes the interatomic vectors . (ii) The outputs and are node properties and are therefore predicted via two distinct multi-layer perceptron heads starting from the node representation of PET. It should be noted that the “raw” FlashMD predictions are chosen to be mass-scaled, i.e. and . This ensures an equal-footings treatment of displacements and momenta for atoms of different mass. Further details on the architecture and the training procedure are available in Appendices A and B, respectively.
@@ 4 Results
To demonstrate the capabilities of FlashMD, we trained two types of models: water-specific models trained on a dataset of MD trajectories for liquid water, and general-purpose, universal models trained on MD trajectories of structures sampled from the MAD dataset (see Ref. [39]). All reference MD simulations were performed with the PET-MAD universal MLIP [39]. For both the water-specific and universal cases, we trained separate models targeting different time strides. Full training details are available in App. B. The following subsections will focus on the testing of such models in predicting meaningful physical observables for the corresponding systems. Accuracies and ablation studies are discussed in App. F, and timings are provided in App. G.
@@ 4.1 Liquid water
Liquid water is central to many physical, chemical, biological and environmental processes, in great part thanks to its microscopic hydrogen-bonding network and resulting physical and chemical properties. The study of liquid water at the microscopic level with MD is therefore a very active area of research [67, 68, 69, 70], and we use it here as an example of how FlashMD models can accurately predict physical observables at the molecular level. For consistency with the universal model, we train the water-specific model on trajectories obtained with PET-MAD, even though its reference electronic-structure method (PBEsol [71]) is known to grossly overestimate the melting point of water. For this reason, we perform simulations with a target temperature of 450 K, above the melting temperature of this model. Results for the q-TIP4P/f model [72] at room temperature are discussed in Appendix I.
We focus on the evaluation of static observables, i.e. the equilibrium, time-independent properties of a physical system which can be estimated as averages over MD trajectories. As easy-to-compute diagnostics, we investigate the mean kinetic energy and atomic radial distribution functions in simulations, as well as the equilibrium density predicted by simulations. The mean kinetic energies (expressed as effective temperatures) are shown in Table 1. It can be seen that, while the models without energy conservation can show pathological deviations from the target temperatures, the energy conservation enforcement approach described in App. C always recovers the correct temperatures for the overall simulation. However, significant deviations can still be observed for the global stochastic velocity rescaling [30] (SVR) thermostat in the kinetic energies resolved by atom type. This is a spurious effect (classically, each degree of freedom should have a mean kinetic energy equal to ) which is also observed in non-conservative force models [21]. The link to direct force prediction (that can be seen as the of trajectory prediction) is also suggested by the fact that lack of equipartition is observed also for short-time FlashMD models, that have excellent validation accuracy. As a consequence, one needs to employ local thermostats, such as those based on Langevin dynamics, similar to what was done in Bigi et al. [21]. With a judicious choice, one can achieve accurate sampling of equilibrium properties, without reducing substantially sampling efficiency. However, local thermostats disrupt dynamical properties, to an extent that depends on the strength of the thermostat coupling, and a thorough quantitative analysis of dynamical properties require disentangling the effect of the long-stride sampling and that of the thermostats needed to obtain accurate equilibrium sampling (see App. I). For these reasons, we primarily focus our quantitative analysis on time-independent equilibrium properties, and discuss examples where FlashMD qualitatively captures time-dependent behavior.
@@ 4.2 Universal long-stride sampling
We extend upon the previous liquid water simulations, generating constant-pressure trajectories of a single alanine dipetide molecule in solution, following closely the setup described in Morrone et al. [76]. The energy landscape is probed in three different simulations: MD using the PET-MAD potential with 0.5 fs stride, and universal FlashMD with 8 fs and 16 fs strides. Note that for this system 0.5 fs is a limit above which MD simulations exhibit severe instabilities. The Ramachandran plots (Figure 3a) show the characteristic distribution of molecular conformers in terms of the backbone dihedral angles. This is a model for the backbone flexibility of proteins and demonstrates the ability of FlashMD to recover major features of the Ramachandran plot with far larger (up to a factor of 32) stride.
@@ 5 Discussion
This is not to say that circumventing Hamiltonian dynamics is without problems. We highlight several ways an architecture similar to FlashMD could fail, by breaking some of the fundamental symmetries and conservation laws that are obeyed (at least approximately) by conventional integrators. We show how these shortcomings can be mitigated, e.g., by performing energy rescaling at inference time, or by including thermostats to control systematic drifts in the constant-energy trajectories. Many of these shortcomings are shared by non-conservative force models, which have also become fashionable as a tool to accelerate MD. We would argue that the transformative speed-up afforded by FlashMD makes direct MD trajectory prediction a more promising approach, at least when performing exploratory studies that require simulating long time scales. As shown concretely by challenging examples that simulate with semi-quantitative accuracy three archetypal systems for biochemistry, surface science and energy technologies, FlashMD can already be applied to realistic simulation problems, capturing the essential equilibrium and dynamical processes while accelerating sampling significantly in many cases (App. G).
@@
Helper functions to download universal FlashMD models and to prepare simulations are distributed with the flashmd package, available on PyPi. An example of the use of FlashMD is also available at https://atomistic-cookbook.org/examples/flashmd/flashmd-demo.html.
@@
The authors would like to thank Davide Tisi and Federico Grasselli for help in processing LiPS trajectories. FB and MC acknowledge support from the NCCR MARVEL, funded by the Swiss National Science Foundation (SNSF, grant number 182892) and from the Swiss Platform for Advanced Scientific Com- puting (PASC). MC and SC acknowledge funding from the Swiss National Science Foundation (Project 200020_214879).
@@ A.1 Why predict and : small- and large-time limits
It should be noted that the models for liquid argon in the early work by Zheng et al. [61] effectively predict and . While the additional terms ( and , respectively) reduce the variance of the targets for small time steps and/or very smooth potential energy surfaces (such as Lennard-Jones argon in Ref. [61]), they instead increase it for more complex systems and larger predicted time steps, which are the focus of the present work. This is a consequence of more complex systems generally having smaller position and momentum correlation times. As a result, we do not shift the targets by these additional terms.
@@ A.2 Predicting mass-scaled positions and momenta
All models shown in this work are trained on, and therefore predict, mass-scaled displacements and momenta defined as and , respectively, where is the mass of atom . This is aimed at making the scales of the training targets uniform across atoms of potentially very different mass, and it is of fundamental importance for models trained on the whole periodic table. This prevents, for example, the displacement of light atoms or the momenta of heavy atoms from dominating the loss, instead leading to good predictions for all atoms, independent of their mass. At prediction time, a simple scaling using the masses recovers the conventional displacement and momenta.
@@ A.3 Center-of-mass enforcement
Both conditions are enforced within the model to avoid unphysical drift effects during molecular dynamics simulations (Fig. 1). We note that removing the center-of-mass motion entirely would not be correct in the general case, although many MD simulations are performed with this additional constraint.
@@ B.2 Training-time energy conservation
We found that the degree of energy conservation on structures of the validation set correlates well with the quality of the models during molecular dynamics runs. As a result, during training, we choose the best model as the one having the lowest product of three terms, evaluated across the validation set: (i) RMSE on the predictions of , (ii) RMSE on the predictions of , and (iii) RMSE on the energy of the resulting structure when compared to the energy of the target structure.
Indeed, as shown in App. F, models with similar accuracy on positions and momenta can predict MD states with highly varying degrees of energy conservation. In particular, energy misalignment is often observed if the error on the energies is ignored.
Furthermore, we found it useful to compare errors in total energies with familiar metrics such as “chemical accuracy” or the accuracy of the underlying energy model. For all models tested in this work, such comparisons correlate extremely well with the quality of the models in the resulting MD simulations.
@@ B.3 Reference MD trajectory generation
A water structure at experimental density (at 298 K and 1 atm) was equilibrated with PET-MAD (or q-TIP4P/f for the q-TIP4P/f-based water models discussed in Appendix I). Subsequently, two more structures were generated by increasing and decreasing the volume of the cell by 10%, scaling the atomic positions accordingly. For each of the three resulting structures, equilibration runs were performed at all temperatures between 20 and 1000 K, in steps of 20 K, with a time step of 0.5 fs and a duration of 5 ps, using a Langevin thermostat with a characteristic time of 10 fs. Subsequently, each equilibrated structure was used to produce an MD trajectory of 2 ps with a time step of 0.25 fs. Structures for training were extracted from these trajectories every 100 fs, and augmented with their time-reversed version, for a total of 5400 structures.
@@ Appendix G Timings
| Acceleration factor | * | * |
@@
In this appendix, we report the overall timings for the simulations that were performed in this work. When multiple FlashMD models with different time steps were used, we report the largest-stride model that affords qualitatively accurate results. The timings obtained in this way are compiled in Table 2. In comparing these results with the theoretical speed-ups given by , one should keep several issues into considerations: (i) We did not fine tune the time step of the base MD runs, nor attempt a fine-grained grid of acceleration factors for the FlashMD models. This might skew results in either directions: for instance, water can be run stably with 0.5 fs MD, and LiPS could have probably been run stably with a 32 fs FlashMD model. (ii) We did not systematically explore the Pareto frontier of FlashMD models in its architecture setup. In the current version, one evaluation of FlashMD is slower than a PET-MAD energy evaluation (by a factor of roughly 50%) due to hyperparameter differences. (iii) Some of the extensions require additional energy evaluations; for instance energy scaling or the integrator require one or two energy evaluations each. There are obvious optimizations, such as only rescaling energy every few steps, which we did not consider, to keep the analysis clearer for this first demonstration of a universal FlashMD model.
Appendix H Liquid argon benchmark

### v2

@@ Abstract
Molecular dynamics (MD) provides insights into atomic-scale processes by integrating over time the equations that describe the motion of atoms under the action of interatomic forces. Machine learning models have substantially accelerated MD by providing inexpensive predictions of the forces, but they remain constrained to minuscule time integration steps, which are required by the fast time scale of atomic motion. In this work, we propose FlashMD, a method to predict the evolution of positions and momenta over strides that are between one and two orders of magnitude longer than typical MD time steps. We incorporate considerations on the mathematical and physical properties of Hamiltonian dynamics in the architecture, generalize the approach to allow the simulation of any thermodynamic ensemble, and carefully assess the possible failure modes of such a long-stride MD approach. We validate FlashMD’s accuracy in reproducing equilibrium and time‐dependent properties, using both system‐specific and general-purpose models, extending the ability of MD simulation to reach the long time scales needed to model microscopic processes of high scientific and technological relevance.
@@ 2.3 ML modeling of molecular dynamics trajectories
Generative ML techniques have been applied to directly sample the thermodynamically accessible system configurations [48, 49, 50, 51], which is especially useful for biomolecular systems with slow conformational transitions. Such models allow cheap prediction of the thermodynamic properties that conventionally require long MD simulations, but they disregard the time-dependent behavior of the systems and hence are unsuitable for investigating the physicochemical phenomena that can only be explained via the system’s dynamics. More recently, the implicit transfer operator (ITO) has been proposed to extend Boltzmann generators to also recover, with stochastic trajectory, a measure of long-time dynamical processes [52]. Another relevant approach is Timewarp [53], which employs a normalizing flow as a proposal distribution within a Markov chain Monte Carlo scheme targeting the Boltzmann distribution, achieving effective time steps on the order of – fs for molecular systems.
@@
This is the class of methods that most closely resembles our approach. It is distinct from the former two approaches in that no generative approaches or multiple time step information is used. In this case, the ML models take only the positions and momenta of atoms of the system at time step as inputs and deterministically predict their changes at , hence “directly propagating” the dynamics. In MDnet by Zheng et al. [62], the chemical system is described as a graph, with the edge features incorporating both the relative positions and momenta between the atoms. The model then predicts the changes in the positions and momenta for a fixed large time step . Very recently, Thiemann et al. have demonstrated TrajCast [63], an autoregressive equivariant network for direct MD prediction. Their framework has been shown to achieve good accuracy in reproducing trajectories for individual molecular or bulk systems at a specific thermodynamic state point. It is also worth mentioning GICnet [64] and its transferable, transformer-based variant MDtrajNet-1 [65], which learns a function that takes as inputs the initial positions, velocity, and to return the positions at time , the Equivariant Graph Neural Operator [66] (EGNO) approach, which predicts the evolution of the system at multiple times using equivariant temporal convolution in Fourier space, and the Graph Network-based Simulators [67, 68] (GNS), which have been developed for arbitrary particle-based systems without the chemical context.
@@ 3 The FlashMD framework
In this section, we explain the design choices made for “FlashMD”, our proposed approach for the direct learning and prediction of MD trajectories.
3.1 Learning molecular dynamics trajectories with graph neural networks
Given that MD shares with interatomic potentials the properties of -equivariance and the use of atomic geometries as inputs, we propose that FlashMD should be built on top of similar GNN architectures to those that have successfully been used to model machine-learning interatomic potentials (MLIPs). In this work, we choose the Point-Edge Transformer [18] (PET), although any GNN architecture could be used. Compared to the original architecture in Ref. [18], we make two physically motivated modifications to adapt it to the prediction of trajectories: (i) Each node state is also initialized using the particle momentum , encoded via a multi-layer perceptron, in addition to the chemical species of the atom under consideration. The initialization of the edge states remains unchanged, and it includes the interatomic vectors . (ii) The outputs and are node properties and are therefore predicted via two distinct multi-layer perceptron heads starting from the node representation of PET.
It should be noted that the “raw” FlashMD predictions are chosen to be mass-scaled, i.e. and . This ensures a treatment of displacements and momenta on equal footing for atoms of different mass, although a data-driven approach is also possible (App. A). Further details on the architecture and the training procedure are available in Appendices A and B, respectively.
@@ 4 Results
To demonstrate the capabilities of FlashMD, we trained two types of models: water-specific models trained on a dataset of MD trajectories for liquid water, and general-purpose, universal models trained on MD trajectories of structures sampled from the MAD dataset (see Ref. [39]). All reference MD simulations were performed with the PET-MAD universal MLIP [39]. For both the water-specific and universal cases, we trained separate models targeting different time strides. Full training details are available in App. B. The following subsections will focus on the testing of such models in predicting meaningful physical observables for the corresponding systems. Ablation studies are discussed in App. F, and timings are provided in App. G. Test set accuracy benchmarks against MDNet [62] and TrajCast [63] are reported in App. H and App. I, respectively.
@@ 4.1 Liquid water
Liquid water is central to many physical, chemical, biological and environmental processes, in great part thanks to its microscopic hydrogen‐bonding network and resulting physical and chemical properties. The study of liquid water at the microscopic level with MD is therefore a very active area of research [70, 71, 72, 73], and we use it here as an example of how FlashMD models can accurately predict physical observables at the molecular level. For consistency with the universal model, we train the water-specific model on trajectories obtained with PET-MAD, even though its reference electronic-structure method (PBEsol [74]) is known to grossly overestimate the melting point of water. For this reason, we perform simulations with a target temperature of 450 K, above the melting temperature of this model. Results for the q-TIP4P/f model [75] at room temperature are discussed in Appendix J.
We focus on the evaluation of static observables, i.e. the equilibrium, time-independent properties of a physical system which can be estimated as averages over MD trajectories. As easy-to-compute diagnostics, we investigate the mean kinetic energy and atomic radial distribution functions in simulations, as well as the equilibrium density predicted by simulations. The mean kinetic energies (expressed as effective temperatures) are shown in Table 1. It can be seen that, while the models without energy conservation can show pathological deviations from the target temperatures, the energy conservation enforcement approach described in App. C always recovers the correct temperatures for the overall simulation. However, significant deviations can still be observed for the global stochastic velocity rescaling [30] (SVR) thermostat in the kinetic energies resolved by atom type. This is a spurious effect (classically, each degree of freedom should have a mean kinetic energy equal to ) which is also observed in non-conservative force models [21]. The link to direct force prediction (that can be seen as the of trajectory prediction) is also suggested by the fact that lack of equipartition is observed also for short-time FlashMD models, that have excellent validation accuracy. As a consequence, one needs to employ local thermostats, such as those based on Langevin dynamics, similar to what was done in Bigi et al. [21]. With a judicious choice, one can achieve accurate sampling of equilibrium properties, without reducing substantially sampling efficiency. However, local thermostats disrupt dynamical properties, to an extent that depends on the strength of the thermostat coupling, and a thorough quantitative analysis of dynamical properties require disentangling the effect of the long-stride sampling and that of the thermostats needed to obtain accurate equilibrium sampling (see App. J). For these reasons, we primarily focus our quantitative analysis on time-independent equilibrium properties, and discuss examples where FlashMD qualitatively captures time-dependent behavior.
@@ 4.2 Universal long-stride sampling
We extend upon the previous liquid water simulations, generating constant-pressure trajectories of a single alanine dipetide molecule in solution, following closely the setup described in Morrone et al. [79]. The energy landscape is probed in three different simulations: MD using the PET-MAD potential with 0.5 fs stride, and universal FlashMD with 8 fs and 16 fs strides. Note that for this system 0.5 fs is a limit above which MD simulations exhibit severe instabilities. The Ramachandran plots (Figure 3a) show the characteristic distribution of molecular conformers in terms of the backbone dihedral angles. This is a model for the backbone flexibility of proteins and demonstrates the ability of FlashMD to recover major features of the Ramachandran plot (particularly the low-energy basins in the range) with strides up to 32 times larger.
@@ 5 Discussion
This is not to say that circumventing Hamiltonian dynamics is without problems. We highlight several ways an architecture similar to FlashMD could fail, by breaking some of the fundamental symmetries and conservation laws that are obeyed (at least approximately) by conventional integrators. We show how these shortcomings can be mitigated, e.g., by performing energy rescaling at inference time, or by including thermostats to control systematic drifts in the constant-energy trajectories. While the design choices of FlashMD deviate from Hamiltonian dynamics, an alternative framework that preserves symplecticity (and therefore Hamiltonian dynamics) in the direct learning of trajectories is proposed elsewhere [81]. Many of the shortcomings of non-symplectic FlashMD are shared by non-conservative force models, which have also become fashionable as a tool to accelerate MD. We argue that the transformative speed-up afforded by FlashMD makes direct MD trajectory prediction a more promising approach, at least when performing exploratory studies that require simulating long time scales. As shown concretely by challenging examples that simulate with semi-quantitative accuracy three archetypal systems for biochemistry, surface science and energy technologies, FlashMD can already be applied to realistic simulation problems, capturing the essential equilibrium and dynamical processes while accelerating sampling significantly in all cases (App. G).
@@
The FlashMD models presented in this work were trained with the metatrain package [82], and they support inference in multiple simulation engines (including ASE [83], i-PI [84], LAMMPS [85]) through metatomic [82].
Helper functions to download universal FlashMD models and to prepare simulations are distributed with the flashmd package available on PyPI. Further information and instructions can be found at https://flashmd.org, including links to the training datasets and scripts to reproduce the reported results on HuggingFace and Materials Cloud [86]. An example of the use of FlashMD is also available at https://atomistic-cookbook.org/examples/flashmd/flashmd-demo.html.
Besides the universal models presented in this work, which were trained to reproduce molecular dynamics at the PBEsol level of theory (through the PET-MAD universal potential), we also make available FlashMD models based on the more accurate r2SCAN functional which we recommend for scientific use.
@@
The authors would like to thank Davide Tisi and Federico Grasselli for help in processing LiPS trajectories. FB and MC acknowledge support from the NCCR MARVEL, funded by the Swiss National Science Foundation (SNSF, grant number 182892) and from the Swiss Platform for Advanced Scientific Computing (PASC). MC and SC acknowledge funding from the Swiss National Science Foundation (Project 200020_214879).
@@ A.1 Why predict and : small- and large-time limits
It should be noted that the models for liquid argon in the early work by Zheng et al. [62] effectively predict and . While the additional terms with respect to FlashMD ( and , respectively) reduce the variance of the targets for small time steps and/or very smooth potential energy surfaces (such as Lennard-Jones argon in Ref. [62]), they instead increase it for more complex systems and larger predicted time steps, which are the focus of the present work. This is a consequence of more complex systems generally having smaller position and momentum correlation times. As a result, we do not shift the targets by these additional terms in our work.
@@ A.2 Predicting mass-scaled positions and momenta
All models shown in this work are trained on, and therefore predict, mass-scaled displacements and momenta defined as and , respectively, where is the mass of atom . This is aimed at making the scales of the training targets uniform across atoms of potentially very different mass, and it is of fundamental importance for models trained on the whole periodic table. This prevents, for example, the displacement of light atoms or the momenta of heavy atoms from dominating the loss, instead leading to good predictions for all atoms, independent of their mass. At prediction time, a simple scaling using the masses recovers the conventional displacement and momenta. We found that a similar, but data-driven, approach can provide the same benefits. This consists of using different standardization factors for different chemical elements, so that training displacements and momenta are scaled to unit standard deviation during training, for each chemical element (i.e., two scaling factors are used for every single chemical element in the dataset: one for displacements, one for momenta).
@@ A.3 Center-of-mass enforcement
Both conditions are enforced within the model to avoid unphysical drift effects during molecular dynamics simulations (Fig. 1). We note that removing the center-of-mass motion entirely would not be correct in the general case, although many MD simulations are performed with this additional constraint. Although we enforced these contraints within the model in this work, we recommend applying them at inference time only in order not to break the locality assumption that underpins our approach.
@@ B.2 Training-time energy conservation
We found that the degree of energy conservation on structures of the validation set correlates well with the quality of the models during molecular dynamics runs. As a result, during training, we choose the best model as the one having the lowest product of three terms, evaluated across the validation set: (i) RMSE on the predictions of , (ii) RMSE on the predictions of , and (iii) RMSE on the energy of the resulting structure when compared to the energy of the target structure. While an energy term might also be included in the loss function, we found that it slows down training significantly (due to evaluations of the energy model and its gradients), without improving the quality of the FlashMD models in any measurable way.
Indeed, as shown in App. F, models with similar accuracy on positions and momenta can predict MD states with highly varying degrees of energy conservation. In particular, energy misalignment is often observed if the error on the energies is ignored. Although we found this approach to improve the quality of the simulations afforded by our water models (both PET-MAD and q-TIP4P/f), we found its impact on universal models to be less dramatic.
We also found it useful to compare errors in total energies with familiar metrics such as “chemical accuracy” or the accuracy of the underlying energy model. For all models tested in this work, such comparisons correlate extremely well with the quality of the models in the resulting MD simulations.
@@ B.3 Reference MD trajectory generation
A water structure at experimental density (at 298 K and 1 atm) was equilibrated with PET-MAD (or q-TIP4P/f for the q-TIP4P/f-based water models discussed in Appendix J). Subsequently, two more structures were generated by increasing and decreasing the volume of the cell by 10%, scaling the atomic positions accordingly. For each of the three resulting structures, equilibration runs were performed at all temperatures between 20 and 1000 K, in steps of 20 K, with a time step of 0.5 fs and a duration of 5 ps, using a Langevin thermostat with a characteristic time of 10 fs. Subsequently, each equilibrated structure was used to produce an MD trajectory of 2 ps with a time step of 0.25 fs. Structures for training were extracted from these trajectories every 100 fs, and augmented with their time-reversed version, for a total of 5400 structures.
@@ Appendix G Timings
| Acceleration factor | * | * | ||
| Acceleration factor (no ECE) | * | * |
@@
In this appendix, we report the overall timings for the simulations that were performed in this work. When multiple FlashMD models with different time steps were used, we report the largest-stride model that affords qualitatively accurate results. The timings obtained in this way are compiled in Table 2. In comparing these results with the theoretical speed-ups given by , one should keep several issues into considerations: (i) We did not fine tune the time step of the base MD runs, nor attempt a fine-grained grid of acceleration factors for the FlashMD models. This might skew results in either directions: for instance, water can be run stably with 0.5 fs MD, and LiPS could have probably been run stably with a 32 fs FlashMD model. (ii) We did not systematically explore the Pareto frontier of FlashMD models in its architecture setup. However, in the current version, one evaluation of FlashMD is faster than a MLIP energy evaluation due to hyperparameter differences. (iii) Some of the extensions require additional energy evaluations; for instance energy scaling or the integrator require one or two energy evaluations each. There are obvious optimizations, such as only rescaling energy every few steps, which we did not consider in order to keep the analysis clearer for this first demonstration of a universal FlashMD model.
Appendix H Liquid argon benchmark (MDNet)